OpenStudy (anonymous):

Can someone helpp? Enter the equation of the line meeting the given conditions. Please put the equation in standard form. Containing A(5, 3) and perpendicular to a line with slope of -2 I got y=2x+9 and its marking it wrong

OpenStudy (anonymous):

You want a slope that is the negative reciprocal of -2 because you want it to be perpendicular. So, you want a slope of +1/2. You have to use the point-slope form for the equation of a line first, before getting the standard form, so you use the point-slope form:\[y - y _{1} = m(x - x _{1})\]where "m" is your slope and (x1, y1) is your point. y - 3 = (1/2)(x - 5) Now, you can convert to standard form: x - 2y = -1

OpenStudy (anonymous):

You can also put it as: 2y - x = 1

OpenStudy (anonymous):

If you want the slope-intercept form: y = (1/2)x + 1/2

OpenStudy (anonymous):

All good now @baraha ?

OpenStudy (anonymous):

Yeah but which one is standard form? because its an online course and i put in 2y-x=1 and it gave me partial credit

OpenStudy (anonymous):

What are your choices? Some authors use the standard form a little differently.

OpenStudy (anonymous):

If you list your choices here, I can help you further.

OpenStudy (anonymous):

no it doesnt give me a blank

OpenStudy (anonymous):

You might have a choice that looks like: x - 2y + 1 = 0 That's another way to go, but if it doesn't like the form 2y - x = 1 then I would use: x - 2y = -1

OpenStudy (anonymous):

A lot of times, the answer demands that the "x" term be first. Other authors will minimize the "-" signs. It's really all the same. Just go with x - 2y = -1 which was the first one I gave you.

OpenStudy (anonymous):

Does it like that one better? @baraha

OpenStudy (anonymous):

almosttt its not giving me the last point lol

OpenStudy (anonymous):

Are you sure it wants the standard form or the slope-intercept form instead?

OpenStudy (anonymous):

standard form

OpenStudy (anonymous):

Then that online course is a little messed up because my answer is correct. I would use x - 2y = -1 but again, the 2 ways are equivalent for standard form.

OpenStudy (anonymous):

yeah it really is... i was also stuck on another problem and gave me partial credit Indicate the equation of the line meeting the given conditions. Please put the equation in standard form. Containing A(1, 3) and B(0, 2)

OpenStudy (anonymous):

That course is really picky. The only other form you could try is: x - 2y + 1 = 0

OpenStudy (anonymous):

 thats what i got